### v1

@@ 4.2 Comparison with NFTs and ERC-20s
In this section, we discuss the distinctions between inscriptions and established Ethereum Request for Comment (ERC) standards such as ERC-20 [10] and ERC-721 [11] concerning tokens and NFTs, respectively.
Comparison with NFTs.  In the Ethereum blockchain, non-fungible tokens (NFTs) [35] are created through smart contracts. Each user receives a unique token ID, affirming their ownership of a specific asset. These assets, which can be JPEG files or CryptoPunk images, are stored off-chain on a server or on the InterPlanetary File System (IPFS) [17].
In contrast to NFTs, inscription-based tokens do not rely on any smart-contract and, thus, do not allow upgrades. The link to the asset file is inscribed into the transaction data. In a blockchain with a maximal native token supply, such as Bitcoin, the amount of inscription is bound by the blockchain network limits. In contrast, NFTs based on smart-contracts lack such limitations, theoretically allowing for unlimited minting. Also, each inscription is allocated a position in the blockchain, which creates the opportunity to derive the additional value of the inscription from the location within the block.
Comparison with ERC-20.  The other application of inscriptions is to create ERC-20-style tokens. As inscription-based tokens are not reliant on smart-contracts, they are not susceptible to the risk of smart-contract upgrades, and their maximal circulating supply is declared once in the deploy operation. The minting and transfer of each new inscription-based token can be tracked directly on the blockchain and cannot be altered.
However, due to their lack of smart-contract support, inscription-based tokens offer limited functionality. They are primarily utilized for speculative purposes and fall into the category of meme-coins [36]. Based on the inscription technology, it is possible to mint a single NFT-style token or a group of tokens with a predefined token supply in circulation. Nevertheless, it is the off-chain indexer, managed by the token’s creator, responsible for ensuring the data integrity of minted tokens. Trading of inscription-based tokens predominantly occurs within inscriptions marketplaces since they are not compatible with the ERC-20 standard required by decentralized exchanges [40].
@@ 5 Data Collection
Table 2: Data set used to analyze inscription events on five chains.
@@
In our study, we analyze the inscriptions recorded on Ethereum and its major EVM-compatible rollups: zkSync Era, Arbitrum, Optimism, and Base. From these L2s, zkSync Era is a ZK rollup, whereas other chains are optimistic rollups.
To achieve this, we collected blockchain data from various sources. For zkSync, sd use data obtained from an archive node and a Web3-compatible API for zkSync Era. For Arbitrum, Base, Ethereum, and Optimism, we use data sources from Nansen BigQuery [24]. These data sources enable us to collect comprehensive data containing all information recorded on evaluated chains. This includes data about blocks, transactions, and events (or logs triggered during the execution of smart-contracts) specifically related to inscriptions. Table 2 provides an overview of the specifics of our dataset.
@@
To identify these inscriptions on-chain, we specifically search for transactions with input call data starting with “0x646174613a”, representing “data:” in ASCII. This criterion establishes crucial conditions for recognizing inscriptions within the blockchain. The blockchains with the most number of transaction inscriptions in our dataset were zkSync Era and Arbitrum. In total, we found $17054466$ transactions containing inscriptions in our zkSync Era data set issued by $481687$ addresses and added to $2809054$ blocks. Similarly, we identified $16309035$ transactions containing inscriptions in Arbitrum, issued by $118544$ addresses and added to $3575299$ blocks. The average number of inscriptions per block are $6.07$ and $4.56$ for zkSync Era and Arbitrum, respectively. We provide similar statistics for Base, Ethereum, and Optimism chains in Table 2. This dataset forms the basis for our empirical analysis of inscriptions.
@@ 6 Empirical Analysis
This section presents the empirical analysis of the inscriptions recorded on Ethereum and its major EVM-compatible rollups: Arbitrum, Base, Optimism, and zkSync Era. We characterize the inscription protocols, operations, tokens, and their trading dynamics. Subsequently, we investigate the impact of sudden surges in inscription transactions on the blockchains’ performance.
@@
Figure 2: Daily distribution of inscription transactions in our dataset on Ethereum and its rollups (Arbitrum, Base, Optimism, and zkSync Era): (a) Absolute number of inscription transactions; and (b) Percentage of inscription transactions compared to the total number of transactions.
@@ 6.2 Inscriptions Characterization
Figure 3: Percentage breakdown of the top 5 inscriptions attributes—protocol, operations, and tokens—for the chains analyzed in our study: Arbitrum, Base, Ethereum, Optimism and zkSync Era.
@@
Our research indicates that inscription tokens are uniquely unified by a protocol-tick pair. Therefore, we aggregate data using this pair to identify the top tokens. Figure 3 illustrates the breakdown of inscription transactions based on protocol, operation, and the leading token. Below, we provide our observations for each of these categories.
@@
Protocols.  Our initial observation reveals that each blockchain contains a particular leading inscription protocol, similar to BRC-20 on Bitcoin, constituting between 54% and 93% of transactions. Moreover, blockchain-specific standards such as FAIR-20 for Arbitrum, ERC-20 for Ethereum, and ZRC-20 for zkSync Era exhibit prominence. The LAYER2-20 protocol facilitates minting and trading of the inscription token across multiple rollups. It is present across all studied L2s, emerging as the dominant inscription protocol on Base and Optimism, and the second largest on zkSync Era.
@@
Operations.  The new inscription standards resemble the original BRC-20 standard but introduce additional operation types such as claim, list, buy, and sell, which were absent in the original BRC-20 implementation on Bitcoin. Notably, the claim operation works similarly to mint, generating new tokens. Hence, we aggregate mint and claim operations, constituting approximately 99% of all transactions related to inscriptions across all studied chains. Understandably, a small number of transactions are attributed to the list operation, used for listing inscriptions for sale in an inscription marketplace, and the deploy operation, which declares the token on the chain along with its token supply. The low percentage of transfer, sell, and buy transactions suggests that inscription tokens are not yet actively traded on EVM chains.
@@
Tokens.  Based on our observations, various protocols deploy and mint tokens with the same ticks. Therefore, we identify each inscription token by a protocol-tick pair. We found that Arbitrum, Base, and Optimism each have one dominant token, representing between 73% and 85% of inscription transactions. The distribution of tokens is most diverse on Ethereum, with ERC-20 NODES token contributing 32% of transactions, and each other tokens less than 3.5%. The four largest tokens on zkSync Era account for over 90% of inscription transactions and zkSync Era has more powerful leading tokens than in other chains. The token LAYER2-20 \$L2 is present on all rollups, being also the dominant token on Base and Optimism.
6.3 Inscription Trading.
@@
In this study, we analyze the trading behavior of two inscription tokens within the zkSync Era: zrc-20 sync, the predominant token, and era-20 bgnt, a token of smaller stature but notable for its completed minting process. Our analysis, presented in Figures 4 and 5, shows their respective trading activities. Notably, era-20 bgnt demonstrates a higher level of trading activity compared to zrc-20 sync. Specifically, we observed $477$ transfer transactions for zrc-20 sync, alongside corresponding token median prices as shown in Figure 4, and $1148$ buy transactions for era-20 bgnt, with corresponding median prices depicted in Figure 5
@@
Figure 4: Price per token at every transaction considering the zrc-20 sync token
Figure 5: Price per token at every buy transaction considering the era-20 bgnt token
@@
During the studied period, the price of zrc-20 sync oscillated between \$0.10 and \$0.4, while era-20 bgnt ranged from \$0.10 to \$0.8. Trading of the era-20 bgnt token continued from its minting throughout the entire studied period. As the trades occur in batches of 4 or 5 tokens, they provide minimal compensation to the seller to cover the gas fees necessary to mint and list the inscriptions. The trading of inscriptions at such low prices is feasible due to the low gas fees on rollups, which ranged at zkSync Era from \$0.05 to \$0.25 during the period (see Figure 6). Next, we examine how spikes in inscription transactions influence gas fees across the network.
@@
Table 4 presents a breakdown of fees incurred during operations on inscriptions. In the first transaction, deploy, the maximum number of tokens that can be minted or claimed in a single blockchain transaction is specified. Therefore, mint and claim operations generally deal with this maximum amount specified in deploy, typically 4 or 5 tokens for mint while sometimes even 1000 for claim, as depicted in Table 1. We calculate the average fee per single token for each operation. We found that the mint operation is more efficient than claim, although both have a similar meaning. Minting was specified in the original BRC-20 specification. The interesting comparison provides claim operation for layer2-20 \$L2 token, as it is operating across all rollups. Its mining (claiming) operation had a lower median fee on Optimism and Base, followed by zkSync Era and Arbitrum.
@@ 6.4 Impact on Gas Fees
Figure 6: Daily transaction count and median transaction cost paid by the user: (a) Transactions count; and (b) Median transaction cost.
@@
(b) Ethereum mainnet, a L1 blockchain.
Figure 7: Distribution of gas fees in GWei relative to transaction amount: (a) L2 Rollups; and (b) Ethereum mainnet.
This section analyzes the impact of daily transaction volumes on the median gas fee per transaction. Unlike L1 blockchains, rollups are expected to offer lower gas fees with increasing transaction counts due to their mechanism. As more L2 transactions can be compressed into a single batch stored in the L1 network, gas fees tend to decrease. Notably, spikes in daily transaction counts on specific days resulted in a decrease in median gas fees per transaction for other users, as illustrated in Figure 6. Fees on L2 rollups are linked to compression and are also affected by the fees of L1 Ethereum.
Further analysis of the relationship between transaction count and median gas fee is presented in Figures 7. Generally, gas fees with more transactions on rollups either remain unchanged (Base, Optimism) or decrease (Arbitrum, zkSync). Conversely, on Ethereum (L1 chains), gas fees tend to increase with higher transaction counts. Additionally, zkSync demonstrates higher efficiency in reducing gas fees compared to Arbitrum, attributable to its ZK rollup design, while Arbitrum, like Base and Optimism, is an optimistic rollup. An inherent characteristic specific to ZK rollups, such as zkSync Era, is their ability to distribute the cost of generating validity proofs across many transactions. ZK rollups utilize cryptographic methods like ZK-STARKs (Zero-Knowledge Scalable Transparent Argument of Knowledge) to validate transactions off-chain and submit them in batches to underlying L1 blockchains. With more transactions in a batch, the cost per transaction decreases. Thus, unlike L1 blockchains, gas costs on zk-rollups decrease with a larger transaction volume, as observed on November 17, December 5, December 17, and December 20 on zkSync Era. Additionally, on March 13, Ethereum’s Deucun upgrade [2] introduced blobs and danksharding aimed at enhancing Ethereum rollups’ efficiency more.
@@ 7 Discussion
In this section, we discuss various aspects of inscriptions, starting with the impact of blobs on inscription ownership. We then explore whether inscriptions are beneficial or detrimental to the blockchain ecosystem. Finally, we address the potential for pump and dump schemes associated with inscriptions.
@@
Impact of blobs on inscription ownership.  On March 13, 2024, Ethereum underwent the Dencun upgrade [2], which introduced temporary data storage called blobs. Blobs, initially designed to increase the efficiency of Ethereum rollups, were used to issue new inscription directly in Ethereum blockchains, called BlobScriptions [1]. Yet, blob data is intended to be stored by Ethereum nodes for 18 days, during which rollup operators or other parties verify their transactions. After the 18-day period, BlobScriptions disappear from the Ethereum blockchain and are only stored (and available) in the indexer of the creator protocol, outside the blockchain. Similarly, rollups are not required to store (outside of blobs) the transaction input data used by rollup-based inscriptions, raising questions about the ownership of inscriptions.
@@
Inscriptions: good or bad?  A report by Binance on BRC-20 highlights that the unexpected surge in transactions resulted in an increase in transaction fees for many blockchains and rollup [32]. The report suggests that such an increase in fees can be seen as a natural progression in blockchain adoption. To illustrate, if a few million more individuals had chosen to use Bitcoin or other blockchains for peer-to-peer transactions, a similar spike in transaction fees would have occurred. Therefore, attribution of this spike solely to BRC-20s is considered irrelevant. What is perhaps more troubling, other protocols, such as Arbitrum, experienced downtime of approximately 78 minutes [33]. Consequently, this led to users worrying about long transaction inclusion times. This led the blockchain community to debate whether ordinals and inscriptions benefit the blockchain users.
However, others might argue that the existing blockchain protocols were not fully equipped to handle such a significant influx of transactions. As described earlier, zkSync Era encountered higher transaction-per-second (TPS) rate, approximately 96% of all transactions were related to inscriptions for a couple of hours, and a reduction in gas fees.
@@
Challenges and opportunities for trading inscription.  Inscriptions are not compatible with ERC-20 tokens and therefore cannot be traded on traditional decentralized exchanges based on automated market makers (AMM) such as Uniswap [3], which is the base for trading tokens with low market capitalization. Today, inscriptions are listed at NFT-marketplaces in a limited-order book type, reducing their liquidity [16, 8, 44]. Also, for many inscription-based meme-coins, their minting processes have not yet finished, and new tokens are expected to be minted until the token limit declared in the deploy operation is fully researched. The wrapper of inscription-based meme-coins into the ERC-20 token could allow one to list of inscriptions at decentralized exchanges and to attract traders of meme-coins that are not used to NFT marketplaces. We see this as the next evolution of inscriptions if they are intended to stay.
Hope or Hype?  During minting periods, platforms utilized various channels to engage users, including communities, influencers, and specific chains that amplified them. Consequently, numerous users issued inscriptions across different platforms [8, 16, 44].
Understanding user expectations, often influenced by FOMO, regarding inscriptions is crucial. For instance, are inscriptions more about hope or hype? This question relates to whether inscriptions are a novel concept or just a passing trend. In the case of Bitcoin, there was a significant surge in interest and user participation during the BRC-20 movement, followed by a subsequent decline in enthusiasm over the following months [36, 37]. Our work shows similar results for EVM-compatible chains: inscriptions peak initially but witness a decline in user interest over time, as shown in Section 6.1.
@@
The evolution of inscriptions fuels discussions on innovative features and integration with tokens and NFTs. Despite potential profitability, marketplace platforms encounter challenges in translating inscriptions into substantial profits for users, particularly across different chains. Due to the nature of inscriptions, which rely on off-chain applications, there is currently no mechanism to prevent users from purchasing previously sold inscriptions. Additionally, users can utilize software programs to automatically claim the majority of inscriptions during the minting or claiming period. Hence, should inscriptions be distributed directly to wallets, similar to traditional airdrop methods [22]?

### v2

@@ 3.2 Comparison with NFTs and ERC-20s
In this section, we discuss the distinctions between inscriptions and established Ethereum Request for Comments (ERC) standards such as ERC-20 [11] and ERC-721 [12] concerning tokens and Non-Fungible Token (NFT), respectively.
Comparison with NFTs.  In the Ethereum blockchain, NFTs [33, 38] are created through smart contracts. Each user receives a unique token ID, affirming their ownership of a specific asset. These assets, which can be JPEG files or CryptoPunk images [33], are stored off-chain on a server or on the InterPlanetary File System (IPFS) [18].
In contrast to NFTs, inscription-based tokens do not rely on any smart contract and, thus, do not allow upgrades. The link to the asset file is inscribed into the transaction data. In a blockchain with a maximal native token supply, such as Bitcoin, the amount of inscription is bound by the blockchain network limits. In contrast, NFTsbased on smart contracts are free from such limitations, theoretically allowing for unlimited minting. Also, each inscription is allocated a position in the blockchain, which creates the opportunity to derive the additional value of the inscription from the location within the block.
Comparison with ERC-20.  The other application of inscriptions is to create ERC-20-style tokens. As inscription-based tokens are not reliant on smart contracts, they are not susceptible to the risk of smart contract upgrades, and their maximal circulating supply is declared once in the deploy operation. The minting and transfer of each new inscription-based token can be tracked directly on the blockchain and cannot be altered.
However, due to their lack of smart contract support, inscription-based tokens offer limited functionality. They are primarily utilized for speculative purposes and fall into the category of meme-coins [39]. Based on the inscriptions technology, it is possible to mint a single NFT-style token or a group of tokens with a predefined token supply in circulation. Nevertheless, it is the off-chain indexer, managed by the token’s creator, responsible for ensuring the data integrity of minted tokens. Trading of inscription-based tokens predominantly occurs within inscriptions marketplaces since they are not compatible with the ERC-20 standard required by DEXs [44].
@@ 4 Data Collection
Table 2: Data set used to analyze inscription events on five blockchains.
@@
We analyze the inscriptions data recorded on Ethereum and its major EVM-compatible rollups: ZKsync Era, Arbitrum, Optimism, and Base. From these L2s, ZKsync Era is a ZKrollup, whereas other chains are optimistic rollups.
We then collected blockchain data from various sources. For ZKsync, we use data obtained from an archive node and a Web3-compatible API for ZKsync Era. For Arbitrum, Base, Ethereum, and Optimism, we use data sources from Nansen BigQuery [26]. These data sources enable us to collect comprehensive data containing all information recorded on the evaluated chains. This includes data about blocks, transactions, and events (or logs triggered during the execution of smart contracts) specifically related to inscriptions. Table 2 provides an overview of the specifics of our dataset.
@@
To identify these inscriptions, we specifically search for transactions with input call data starting with “0x646174613a”, representing “data:” in ASCII. This criterion allows us to recognize inscriptions within the blockchain. The blockchains with the most number of transaction inscriptions in our dataset were ZKsync Era and Arbitrum. In total, we found $17054466$ transactions containing inscriptions in our ZKsync Era data set issued by $481687$ addresses and added to $2809054$ blocks. Similarly, we identified $16309035$ transactions containing inscriptions in Arbitrum, issued by $118544$ addresses and added to $3575299$ blocks. The average number of inscriptions per block are $6.07$ and $4.56$ for ZKsync Era and Arbitrum, respectively. We provide similar statistics for Base, Ethereum, and Optimism chains in Table 2. This dataset forms the basis for our empirical analysis of inscriptions.
@@ 5 Empirical Analysis
In this section, we present our empirical analysis of the inscriptions recorded on Ethereum and its major EVM-compatible rollups: Arbitrum, Base, Optimism, and ZKsync Era. We characterize the inscription protocols, operations, tokens, and their trading dynamics. Then, we investigate the impact of sudden surges in inscription transactions on the blockchains’ performance.
@@
Figure 1: Daily distribution of inscription transactions in our data set: (a) absolute number of inscription transactions; and (b) percentage of inscription transactions compared to the total number of transactions.
@@ 5.2 Inscriptions Characterization
Figure 2: Percentage breakdown of the top 3 inscriptions attributes — protocol, operations, and tokens — for the chains analyzed in our study: Arbitrum, Base, Ethereum, Optimism and ZKsync Era.
@@
Our analysis indicates that inscription tokens are uniquely unified by a protocol-tick pair. Therefore, we aggregated data using this pair to identify the top inscriptions tokens. Fig. 2 illustrates the breakdown of inscription transactions based on protocol, operation, and the leading token. In the following, we provide our observations for each of these categories.
@@
Protocols.  Our initial observation reveals, in (Fig. 2 row Top Protocols), that each blockchain contains a particular leading inscription protocol, similar to BRC-20 on Bitcoin [46, 40], constituting between 54% and 93% of transactions. In addition, blockchain-specific standards such as FAIR-20 for Arbitrum, ERC-20 for Ethereum, and ZRC-20 for ZKsync Era appear the most. The LAYER2-20 protocol is present across all studied L2s, being the dominant inscription protocol on Base and Optimism, and the second largest on ZKsync Era.
@@
Operations.  Recall from §3 that both claim and mint operation types work similarly. Thus, if evaluated together they account for approximately 99% of all inscription transactions in all chains we studied. From Fig. 2 row Top Operations, a small number of transactions are attributed to the list operation, used for listing inscriptions for sale in an inscription marketplace, and the deploy operation, which declares the token on the chain along with its token supply. The low percentage of transfer, sell, and buy transactions suggests that inscription tokens are not yet actively traded on EVMchains.
@@
Tokens.  From our analysis (Fig. 2, row Top Tokens), we observe that various protocols deploy and mint tokens with the same tick symbols. To account for this, we identify each token using a protocol-tick pair. Our findings indicate that Arbitrum, Base, and Optimism each have a dominant token, comprising between 74% and 89% of inscription transactions on their respective platforms. In contrast, the token distribution on Ethereum is more diverse, with the ERC-20 NODES token contributing approximately 32% of the transactions, while all other tokens individually represent less than 3.14%. On ZKsync Era, the top four tokens account for over 92% of inscription transactions, also indicating token distribution concentration when compared to other chains. Notably, the token LAYER2-20 \$L2 is present across all rollups and is also the dominant token on both Base and Optimism.
5.3 Inscription Trading
@@
We analyze the trading behavior of two inscription tokens within the ZKsync Era: zrc-20 sync (the predominant token) and era-20 bgnt (a smaller but noteworthy token due to its completed minting process). Our analysis, presented in Fig. 3, shows their respective trading activities. Notably, era-20 bgnt has a higher level of trading activity compared to zrc-20 sync. Specifically, we observed $477$ transfer transactions for zrc-20 sync, along with the corresponding token median prices as shown in Fig. 3(a), and $1148$ buy transactions for era-20 bgnt, with corresponding median prices depicted in Fig. 3(b)
(a) Transfer transactions of zrc-20 sync.
(b) Buy transactions of era-20 bgnt.
@@
Figure 3: Price per token at every transaction on ZKsync Era considering zrc-20 sync and era-20 bgnt tokens.
@@
Our analysis shows that the price of zrc-20 sync oscillated between \$0.10 and \$0.4, while era-20 bgnt ranged from \$0.10 to \$0.8. The trading of the era-20 bgnt token continued from its minting throughout the entire studied period. As the trades occur in batches of 4 or 5 tokens, they provide minimal compensation to the seller to cover the gas fees necessary to mint and list the inscriptions. However, trading inscriptions at such low prices is feasible due to the low gas fees on rollups, which ranged at ZKsync Era from \$0.05 to \$0.25 (see Fig. 4).
@@
Next, we analyze how spikes in inscription transactions influence gas fees across the network. Table 4 presents a breakdown of fees incurred during inscription operations. Recall that, in the first transaction, deploy, the maximum number of tokens that can be minted or claimed in a single transaction is specified. Then, we calculate the average fee per single token for each operation. We found that the mint operation is more efficient than claim, although both have a similar functionality. The interesting comparison provides claim operation for layer2-20 \$L2 token, as it is operating across all rollups. Its mining (claiming) operation had a lower median fee on Optimism and Base, followed by ZKsync Era and Arbitrum.
@@ 5.4 Impact on Gas Fees
Figure 4: Daily transaction count and median transaction cost in USD paid by the user: (a) transactions count; and (b) median transaction cost.
@@
(b) Ethereum mainnet (L1).
Figure 5: Distribution of gas fees in GWei relative to transaction count: (a) L2 rollups; and (b) Ethereum mainnet.
This section analyzes the impact of daily transaction volumes on the median gas fee per transaction. Unlike L1blockchains, rollups are expected to offer lower gas fees with increasing transaction counts due to their scaling mechanism. As more L2transactions can be compressed into a single batch stored in the L1 network, gas fees tend to decrease. In particular, spikes in daily transaction counts on specific days resulted in a decrease in median gas fees per transaction for other users, as shown in Fig. 4. Fees on L2rollups are linked to compression and are also affected by the fees of L1Ethereum.
Further analysis of the relationship between transaction count and median gas fee is presented in Fig. 5. Generally, gas fees with more transactions on rollups either remain unchanged (Base, Optimism) or decrease (Arbitrum, ZKsync). Conversely, on Ethereum, gas fees tend to increase with higher transaction counts. Additionally, ZKsync demonstrates higher efficiency in reducing gas fees compared to Arbitrum, attributable to its ZKrollup design, while Arbitrum, like Base and Optimism, is an optimistic rollup. Thus, unlike L1 blockchains, gas costs on ZKrollups decrease with a larger transaction volume, as observed on November 17, December 5, December 17, and December 20 on ZKsync Era.
@@ 6 Discussion
In this section, we discuss various aspects of inscriptions, starting with the impact of blobs on inscription ownership. We then explore whether inscriptions are beneficial or detrimental to the blockchain ecosystem. Finally, we address the potential for pump-and-dump schemes [21, 43] associated with inscriptions.
@@
Impact of blobs on inscription ownership.  As inscriptions are appended to transactions as arbitrary data within the input call data, there is no assurance of perpetual availability of this data in the future. Blobs, initially designed to increase the efficiency of Ethereum rollups [2], were used to issue new inscription directly in Ethereum blockchains, called BlobScriptions [1]. Yet, blob data is intended to be stored by Ethereum nodes for 18 days, during which rollup operators or other parties verify their transactions. After this period, BlobScriptions disappear from the Ethereum blockchain and are only stored (and available) in the indexer of the creator protocol, outside the blockchain. Similarly, rollups are not required to store (outside of blobs) the transaction input data used by rollup-based inscriptions, raising questions about the ownership of inscriptions.
@@
Inscriptions: good or bad?  A report by Binance on BRC-20 highlights that the unexpected surge in transactions resulted in an increase in transaction fees for many blockchains [35]. The report suggests that such an increase in fees can be seen as a natural progression in blockchain adoption. Therefore, attribution of this spike solely to BRC-20s is considered irrelevant. More concerning, however, is that other protocols, such as Arbitrum, experienced downtime of approximately 78 minutes [36], while ZKsync Era saw an increase in TPS, with nearly 96% of all transactions being related to inscriptions over the course of several hours. Thus, causing users to worry about delayed transaction inclusion. This sparked a broader debate within the blockchain community about whether ordinals and inscriptions truly benefit users.
@@
Challenges and opportunities for trading inscription.  Inscriptions are not compatible with ERC-20 tokens. Therefore cannot be traded on traditional decentralized exchanges based on Automated Market Makers (AMMs)such as Uniswap [3], which is the base for trading tokens with low market capitalization. Today, inscriptions are listed at NFT-marketplaces in a limited-order book type, reducing their liquidity [8, 17, 48]. Also, for many inscription-based meme-coins, their minting process may still be ongoing, and new tokens could continue to be minted until the token limit specified in the deploy operation is reached. Furthermore, wrapping inscription-based meme coins into ERC-20 tokens could allow these inscriptions to be listed on DEXs, attracting meme-coin traders who are not familiar with NFTmarketplaces. We see this as the next evolution of inscriptions if they are intended to stay.
@@
Hope or Hype?  During minting periods, platforms utilized various channels to engage users, including communities, influencers, and specific chains that amplified them. Consequently, numerous users issued inscriptions across different platforms [8, 17, 48]. Understanding user expectations, often influenced by FOMO, regarding inscriptions is crucial. For instance, are inscriptions more about hope or hype? This question relates to whether inscriptions are a novel concept or just a passing trend. In the case of Bitcoin, there was a significant surge in interest and user participation during the BRC-20 movement, followed by a subsequent decline in enthusiasm over the following months [39, 40]. Our work shows similar results for EVM-compatible chains: inscriptions peak initially but witness a decline in user interest over time, as shown in Section 5.1. The evolution of inscriptions fuels discussions on innovative features and integration with tokens and NFTs. Despite potential profitability, marketplace platforms encounter challenges in translating inscriptions into substantial profits for users, particularly across different chains. Due to the nature of inscriptions, which rely on off-chain applications, there is currently no mechanism to prevent users from purchasing previously sold inscriptions. Additionally, users can utilize software programs to automatically claim the majority of inscriptions during the minting or claiming period. Hence, should inscriptions be distributed directly to wallets, similar to traditional airdrop methods [24]?
@@
7 Related Work
Novelty of our approach.  To the best of our knowledge, this is the first work to conduct an in-depth analysis of inscriptions in multi EVM-compatible chains and the user dynamics that arose from the inscription boom observed in the end of 2023 and early 2024. While Bitcoin originally set the stage for this user trend [39, 40], EVMchains have witnessed substantial adoption as well. The absence of analysis in this domain introduces a notable gap in the existing research landscape, which we address in this paper.